For the 2 points, 3,3 and 5,7 generate the equation for a straight line
using slope, y-intercept procedures. Could you do this problem step by step so I can understand how to work this type of problem out. Thanks

Answer by Alan3354(69443)

For the 2 points, 3,3 and 5,7 generate the equation for a straight line
using slope, y-intercept procedures.
----------------
Find the slope, m.
m = diffy/diffx = 4/2
m = 2
-----------
Then use either point in
y-y1 = m*(x-x1) where (x1,y1) is the point
y-3 = 2*(x-3) = 2x-6
y = 2x-3 (slope-intercept form)
2x-y = 3 (standard form)